OpenStudy (anonymous):

Find w(6) for the function w(x) = 4x + 5. (Points : 1) 29 15 44 51

OpenStudy (bibby):

plug in 6. w(6) = 4(6) + 5 = ?

OpenStudy (anonymous):

i dont understand

OpenStudy (the_fizicx99):

4*6 = ?

OpenStudy (anonymous):

24?

OpenStudy (the_fizicx99):

24 + 5 = ?

OpenStudy (anonymous):

29 thanks so much

OpenStudy (the_fizicx99):

Simply replace every x for 6 and simplify

OpenStudy (anonymous):

sweet thanks can u check one for me

OpenStudy (the_fizicx99):

ya

OpenStudy (anonymous):

f (x) = 1/2x − 1 f (x) = −2x − 1 f (x) = 2x − 1 f (x) = −2x + 1

OpenStudy (the_fizicx99):

The y intercept is -1, and the slope is 2.

OpenStudy (anonymous):

so it is c?

OpenStudy (the_fizicx99):

Yes

OpenStudy (anonymous):

or no its B right?

OpenStudy (the_fizicx99):

why would it be B?

OpenStudy (anonymous):

i thought it for a min can u help me with 4 more?

OpenStudy (the_fizicx99):

Oh no wait, it is B, you're moving to the left which is negative :P 2/-1 = -2

OpenStudy (anonymous):

ur positive

OpenStudy (the_fizicx99):

Ya ask @bibby <-- meh bfff ph.D 60 years of experience friend

OpenStudy (bibby):

one problem per question

OpenStudy (anonymous):

f (x) = x f (x) = 1 f (x) = −x f (x) = x + 1

OpenStudy (bibby):

B for the 2nd one. I'm not even looking at the one you just posted

OpenStudy (anonymous):

y not please i need help or ima fail my semester ples

OpenStudy (the_fizicx99):

I was right :> it was B *props to himself* and I gtg, sowwie. Bai.

OpenStudy (anonymous):

i need to graduate im 20 and cant yet till i pass tommorrows the last day

OpenStudy (bibby):

close this question, post another. I'm 32 and am probably going to kill myself in the next 4 years

OpenStudy (the_fizicx99):

.-. wow

OpenStudy (beccaboo333):

LOL